OpenStudy (anonymous):

I have to simplify some radicals... And, I think I'm starting to get the hang of it, but could someone look at my answers to see if they're right or not, and if they aren't give me a mini lesson?

OpenStudy (anonymous):

This is what I have gotten to so far: \[\sqrt{324}=\sqrt{(18)(18)}= \sqrt{18}\times \sqrt{18}\]

OpenStudy (jdoe0001):

looks good

OpenStudy (anonymous):

What's the final answer? @jdoe0001

OpenStudy (jdoe0001):

\(\bf \sqrt{324}\implies \sqrt[2]{324}\implies \sqrt[2]{18\times 18}\implies \Large \sqrt[{\color{red}{ 2}}]{18^{\color{red}{ 2}}}\) well, what do you think we can take out of the radical?

OpenStudy (jdoe0001):

recall, what comes out of the radical, is any factors that match the root

OpenStudy (jdoe0001):

\(\bf \sqrt{324}\implies \sqrt[2]{324}\implies \sqrt[2]{18\times 18}\implies \Large \sqrt[{\color{red}{ 2}}]{18^{\color{red}{ 2}}}\implies 18\)

OpenStudy (anonymous):

So it's 18 then?!

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

since it's exponent matches the root, it comes out

OpenStudy (anonymous):

Thank you! @jdoe0001

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

Could I have you check another answer?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

Is this right?\[6\sqrt{32}=6\sqrt{(8)(4)}=6\sqrt{8}\times \sqrt{4}=6\sqrt{4}\times \sqrt{8}=12\sqrt{8}\]

OpenStudy (jdoe0001):

\(\bf 6\sqrt{32}\implies 6\sqrt{8\cdot 4}\implies 6\sqrt{4}\cdot \sqrt{{\color{blue}{ 8}}}\implies 6\sqrt{2^2}\cdot \sqrt{{\color{blue}{ 2^2\cdot 2}}} \\ \quad \\ 6\cdot 2\cdot 2\sqrt{2}\implies 24\sqrt{2}\)

OpenStudy (anonymous):

That last part confuses me once you highlighted the first 8. @jdoe0001

OpenStudy (jdoe0001):

\(\bf 8\implies 2\cdot 2\cdot 2\implies 2^2\cdot 2\)

OpenStudy (anonymous):

Oh, okay. That part I've got now, but then the last part on the bottom row doesn't make sense! Can you explain what you're doing there?

OpenStudy (jdoe0001):

\(\bf 6\sqrt{32}\implies 6\sqrt{8\cdot 4}\implies 6\sqrt{4}\cdot \sqrt{{\color{blue}{ 8}}}\implies 6\sqrt{2^2}\cdot \sqrt{{\color{blue}{ 2^2\cdot 2}}} \\ \quad \\ 6\sqrt{2^2}\cdot \sqrt{2^2}\cdot \sqrt{2}\implies 6\cdot 2\cdot 2\sqrt{2}\implies 24\sqrt{2}\)

OpenStudy (jdoe0001):

the 2's whose exponent match the root, come out one of the 2's is left behind, since it's not \(\bf 2^2\)

OpenStudy (anonymous):

Okay. That was a tricky one, but I get it now, and the whole idea to do the rest of my work! Thanks again for your time! @jdoe0001

OpenStudy (jdoe0001):

yw